OpenStudy (anonymous):

x+y=7,x+y=-7 solve by elimination

OpenStudy (anonymous):

x+y=7 x+y=-7 There is no solution

OpenStudy (anonymous):

x+y=7 -1(x+y...-7) = -x-y=7 ------------------- x+y=7 -x-y=7 0=7 no solution